OpenStudy (anonymous):

HELP ME PLEASE <3 find to the nearest tenth, the positive value of x in the equation sqrt x^2+21 =2x 1. 7 2. 4.6 3. 2.6 4. 1.3

OpenStudy (anonymous):

Is 21 in the square root or not?

OpenStudy (anonymous):

x=21 I removed the square root and got that for x

OpenStudy (anonymous):

its in square root :)

OpenStudy (anonymous):

Ok. Then square both side first. You get x^2 + 21 = 4x^2. Simplify. 3x^2 = 21 Divide each side by 3. x^2 = 7. Your answer is 2.6 approximately.

OpenStudy (anonymous):

another way: square both sides to get x^2+21 = 4x^2 . after that ,combine like terms 3x^2 = 21 x^2 = 7 after you solve for x, you get x = sqrt(7) . also known as 2.6 :)

OpenStudy (anonymous):

Thank you guys!! xo

OpenStudy (anonymous):

No problem! :)